John Fleming, chief administrator for Valley View Hospital, is concerned about the costs for tests in the hospitals lab. Cha3-a. Compute the variable overhead rate and efficiency variances. 3-b. Is there any relation between the variable overhead efComplete this question by entering your answers in the tabs below. Req 1 Req 2A Req 2B Req ЗА Req 3B For labor cost in the laComplete this question by entering your answers in the tabs below. Req 1 Req 2A Req 2B Req ЗА Req 3B In most hospitals, one-hComplete this question by entering your answers in the tabs below. Req 1 Req 2A Req 2B Req ЗА Req 3B Compute the variable ove
Answer #1

Solution 1:

Standard quantity of material for actual test = (750+3000) * 3 = 11250 plates

Actual quantity of material purchased = 15000 plates

Actual quantity of material used = 15000 - 2000 = 13000 plates

Standard price of material = $3.40

Actual price of material = $48,450 / 15000 = $3.23

Material price variance = (SP - AP) * AQ purchased = ($3.40 - $3.23) * 15000 = $2,550 F

Material quantity variance = (SQ - AQ) * SR = (11250 - 13000) * $3.40 = $5,950 U

Solution 2a:

Standard hours of direct labor = 750*0.8 + 3000*0.4 = 1800 hours

Standard rate of direct labor = $11.20 per hour

Actual hours of direct labor = 2000

Actual rate of direct labor = $21,600 / 2000 = $10.80 per hour

Direct labor rate variance = (SR - AR) * AH = ($11.20 - $10.80) * 2000 = $800 F

Direct labor efficiency variance = (SH - AH) * SR = (1800 - 2000) * $11.20 = $2,240 U

Solution 2b:

As a standard, one-half of the workers in the lab are senior technicians and one-half are assistants. However valley view hospital employs only one-fourth senior technicians and three-fourths assistants. This is resulting in favorable direct labor rate variance as direct labor rate per hour is $10.80 from standard rate of $11.20 per hour. However as hospital employs more assistant, it is resulting in to extra time taken by assistant from standard time because they in experienced resulting in unfavorable labor efficiency variance. Unfavorable labor efficiency variance is higher than favorable direct labor rate variance, therefore this policy should be discontinued.

Solution 3a:

Standard hours of direct labor = 750*0.8 + 3000*0.4 = 1800 hours

Standard rate of variable overhead= $7.10 per hour

Actual hours of direct labor = 2000

Actual rate of variable overhead = $15,200 / 2000 = $7.60 per hour

Variable overhead rate variance = (SR - AR) * AH = ($7.10 - $7.60) * 2000 = $1,000 U

Variable overhead efficiency variance = (SH - AH) * SR = (1800 - 2000) * $7.10 = $1,420 U

Solution 3b:

Yes, there is relation between variable overhead efficiency variance and the labor efficiency variance. Variable overhead is based on direct labor hours, therefore if direct labor efficiency variance improves, variable overhead efficiency variance also improve. Both are positively corelated.